OpenStudy (anonymous):

Solve 4x - 2x - 2 = -42. -20 -7 -5 0

jimthompson5910 (jim_thompson5910):

since 4-2 = 2 this means 4x-2x = 2x

jimthompson5910 (jim_thompson5910):

so 4x - 2x - 2 = -42 turns into 2x - 2 = -42

jimthompson5910 (jim_thompson5910):

What's your next step?

OpenStudy (anonymous):

I have no idea i was hoping someone would explain the steps to find the answer

jimthompson5910 (jim_thompson5910):

your next step is to undo the -2 by adding 2 to both sides

jimthompson5910 (jim_thompson5910):

what does that give you